OpenStudy (anonymous):

the product of th page numbers on two facing pages of a book is 210 find the page numbers

OpenStudy (anonymous):

x(x+1)=210 14,15

OpenStudy (anonymous):

a(a+1)=210 a^2+a-210=0 (a+15)(a-14)=0 a=14 so pages rae 14,15